# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-Modal Consistency in Visual-Stimulated Narratives

### v2

@@ IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and VSNs. This limitation impairs topic representation model, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [41]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [81, 82] (which measure the alignment between text and images through contrastive learning), we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 4.
This network comprises several key components:
@@
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising textual-encoder and visual-encoder to process different modalities. In this work, we adopted the visual-encoder ViT and textual-encoder RoBERTa from the pretrained Chinese-CLIP model, as mentioned in Section IV-B. We adopt this model because it can effectively align the embeddings of images and Cantonese text in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the embeddings $E_{I}$ (from the [CLS] token) and $E_{T}$ (via mean pooling) from the final projection layer of their respective encoders:
$E_{I}=\text{Chinese-CLIP}_{ViT}(I)\in\mathbb{R}^{J\times H}$
$E_{T}=\text{Chinese-CLIP}_{RoBERTa}(T)\in\mathbb{R}^{K\times H}$
where $H$ is the hidden dimension of the extracted embeddings, $J$ is the number of image frames (e.g., $J=1$ for the ADReSS dataset and $J=15$ for the CU-MARVEL-RABBIT corpus), and $K$ is the temporal length of the investigated text sequence.
@@
where [;][;][ ; ] denotes the concatenation operation. This concatenation operation offers at least two advantages. First, it allows the subsequent layers to share the same projection layers for both image and textual embeddings, which is parameter-efficient and feasible since the embeddings are aligned to a common multimodal space during the pre-training stage. Second, inputting the concatenated embeddings into the Attention mechanism enables the model to capture combined information from image-image, image-text, and text-text pairs, thereby enhancing the richness of the contextual features.
@@
As shown in Fig. 4(b), the Attention module aims to capture the temporal dynamics and contextual relationships within and between modalities, by the attending on specific positions of the projected embeddings.
To effectively process dynamic temporal information, we incorporate Rotary Position Embedding (RoPE) [83] with the Attention mechanism. RoPE encodes positional information into token embeddings without disrupting the invariance properties of self-attention, making it well-suited for capturing temporal dependencies. Mathematically, the positional encoding vector for the $i$-th position, $p_{i}\in\mathbb{R}^{2H^{\prime}}$, is defined using sinusoidal functions:
@@
Each element of $p_{i}$ is derived from sinusoidal functions to capture the position information: $\sin(i)$ and $\cos(i)$ are applied to alternate dimensions, with increasing frequencies across dimensions to capture information at various scales. Given the different temporal strides of images and texts, RoPE is applied separately to each modality with different frequencies, i.e., $p_{i}\in\mathbb{R}^{1,\ldots,H^{\prime}}$ for images and $p_{i}\in\mathbb{R}^{H^{\prime}+1,\ldots,2H^{\prime}}$ for texts.
@@
where $Q^{\prime}$ and $K^{\prime}$ are the RoPE-transformed queries and keys. Mask ensures the padding positions of input get near-zero attention weights after the softmax. The scaling factor $\sqrt{H^{\prime}}$ prevents the dot products from growing too large in magnitude. The values remain untransformed to retain the original contents that are being passed through the network.
In VSN tasks such as the Rabbit Story, narrative transcriptions must align with reference images. Participants with NCDs may misidentify animal names (anomia) or miss detailed scenarios (visual perception) in these images. For example, most NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The attention module in the proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
TABLE III: Classification and Regression Results on the CU-MARVEL-RABBIT corpus. Feature dimensions are shown in parentheses. “*”: the baseline systems; $\downarrow$”: lower is better; “Emb.”: embedding(s). Best results are marked in red, and the second-best are underlined.
@@
D5. Prediction
Finally, the up-projected embedding is temporal pooled with Attention mask ($\in\mathbb{R}^{J+K}$) and passed through a fully connected layer to generate the prediction:
$\hat{y}=\sigma(E_{pool}W_{fc}^{T}+b_{fc})\in\mathbb{R}^{C}$
where $C$ is the prediction dimension (number of classes for classification tasks and 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the fully connected layer’s weight and bias, and $\sigma$ is an activation function (softmax for classification tasks and identity for the regression task).
Throughout the network, residual connections are employed at each stage to facilitate gradient flow and mitigate the vanishing gradient problem.
The TITAN model, with the combination of feature extraction, embedding fusion, and Attention-based contextual alignment, can effectively integrate text and image information to evaluate the consistency and temporal flow in narratives. As shown in Fig. 4(c), the heatmap illustrates how the model captures the alignment between text chunks (x-axis) and a sequence of images (y-axis). The varying intensity of the colors represents the degree of alignment between the image and textual content, and the diagonal intensity represents the temporal consistency and flow of the narrative. Bright spots indicate high alignment, showing that text chunks and images correlate strongly at those points, while the diagonal progression reflects the sequence in which events or scenes unfold coherently over time. Moreover, compared to Fig. 3(c, d) using discrete, term-based topics (e.g., limited to 5 coarse topics), Fig. 4(c) leverages instance-level image-text matching. This contrast demonstrates that representing evolving topics with images enables a finer-grained temporal alignment between narratives and their corresponding visual concepts. These heatmaps also provide clinicians a clear view into how textual narratives and visual elements are cohesively aligned to support the model’s evaluation of narrative consistency and temporal flow.
V Experiments & Results
@@ V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and classifier models. Eight systems are designed, each utilizing distinct feature sets. System 1 employs 10 acoustic features, System 2 incorporates 13 linguistic features, and System 3 combines these microstructural features. System 4 utilizes 16 proposed reference-based metrics, and System 5 employs 6 proposed DTM-based features. System 7 combines these macrostructural features, while System 7 integrates all of these microstructural and macrostructural features. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with 5 components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and soft margin $C=1$. The hyper-parameters “components” of PCA, and “kernel” and “soft margin” of SVM are grid searched using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 4, is connected with a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as 5.
@@ V-B Evaluation Protocol
For the evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, we used the F1-score and AUC as the primary metrics, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Despite of AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we normalized the ground-truth labels to the range [0, 1] and used the coefficient of determination ($R^{2}$, ranging from $\infty$- ∞ to 1) and Root Mean Square Error (RMSE, ranging from 0 to +$\infty$+ ∞) as metrics. To ensure robust evaluation, we averaged the results from the final five epochs of each experiment. Notably, both the ADReSS and CU-MARVEL-RABBIT test sets were unseen and maintained an equal distribution of positive and negative samples, ensuring a balanced evaluation.
@@ V-C Results
Table III presents the results of classification and regression tasks, comparing various feature sets and models using F1 as the primary evaluation metric. Among the systems, the proposed TITAN model (Sys. 8), which utilizes combined text and image embeddings, achieves the best performance in both tasks generally, with an F1 score of 0.7238 and AUC of 0.8120 in classification, and R2 of 0.3876 in the regression task. In contrast, linguistic features-based system (Sys. 2) shows an inferior performance among these systems, which is limited by the ASR, segmentation and tagger systems that are not tailored for colloquial Cantonese narratives.
Another finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6947 versus 0.5843 in classification, and R2 of 0.2233 versus 0.1604 in regression, indicating the efficacy of discourse-level analysis on NCD narratives.
@@ VI Analysis & Discussion
Figure 5: Correlation coefficients between proposed features and NCD labels, highlighting the statistical significance of each feature on the x-axis. The amplitudes of the green (macrostructural) and red (microstructural) bars represent correlation values, indicating that most of the top-ranked features are macrostructural.
VI-A Comparison with previous literature
To evaluate the performance of the models, we conducted a cross-comparison study on the public ADReSS dataset. The results in Table IV demonstrates that the proposed TITAN model (Sys.8) achieves superior performance with the F1 score of 0.8889, exceeding others by at least 3.13% relatively.
TABLE IV: Cross comparison study (binary classification) on ADReSS.
Reference F1 AUC Rec. Prec. Acc.
Yuan’2020 [84] 0.8619 - 0.8540 0.8700 0.8540
Zhu’2023 [39] - - - - 0.8224
Botelho’2024 [38] $0$ 0.8330 0.7690 0.8130
Li’2024 [41] 0.8510 - 0.8330 0.8700 0.8540
Ours (Sys.8) 0.8889 0.9410 0.8333 0.9524 0.8958
@@ VI-B Statistical Features Correlation Ranking
To evaluate the significance of the proposed statistical features, we performed Spearman’s rank correlation test to rank all features based on the absolute values of correlation coefficients. As illustrated in Fig. 5, the correlation distribution indicates that the macrostructural features demonstrate greater importance than microstructural features. Notably, DTM-based metric “topic consistency”, acoustic feature “speaking rate”, and several reference-based coverage metrics ranked relatively high with statistical significance ($p<0.001$), suggesting their potential as indicative biomarkers. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features exhibit positive correlations with NCDs. These findings underscore the potential utility of these features for effective classification and a deeper understanding of NCD-related characteristics.
@@ VI-C Interpretable Feature Contribution Analysis
To elucidate the role of the proposed features during the model’s decision pathways, we quantify their contribution using SHapley Additive exPlanations (SHAP) values [85]. SHAP is an interpretable method rooted in cooperative game theory that attributes the prediction of a model to each input feature fairly and consistently using Shapley values. Specifically, for each feature $x_{i}\in X$ (the complete feature set), the SHAP value $\phi_{i}$ represents its average differences in the model’s predictions when the target feature is added to all possible subsets $S$ of other features, i.e.
@@
where $w_{i}=|S|!(|X|-|S|-1)!/|X|!$ denotes the weight accounting for subset permutations, ||$\cdot$|| ⋅ | denotes set cardinality, and $f(S)=ln{\frac{P(\hat{y}|S)}{1-P(\hat{y}|S)}}$ is the log-odds ratio of the model’s output probability $P$ based on given inputs $S$.
Figure 6: Global SHAP value distribution (beeswarm plot) across all test samples in CU-MARVEL-RABBIT. Features are ranked vertically by mean absolute SHAP values (most impactful at top). Each point represents a single instance: horizontal position indicates the SHAP value (negative/positive impact), color reflects the standardized feature value (blue = low, red = high), and point density along each feature shows value distribution.
@@
This formulation enables global feature importance analysis through aggregation of SHAP values across the entire unseen test set. Figure 6 presents the global feature contributions ranked by their absolute SHAP values. Notably, the proposed macrostructural features, including “topic_variability”, “topic_change_rate”, “topic_consistency” and “ROUGE-L”, emerged as the top-ranked global contributors. These findings highlight the critical role of macrostructural patterns in distinguishing NCDs from HCs. The consistent directionality of their SHAP values (positive/negative associations with NCD classification) further validates their reliable discriminative power across diverse instances. For example, the NCD narrative in Fig. 2(b) exhibits low topic variability (original value: 0.25, normalized value: -1.71), high topic change rate (original: 0.61, normalized: 0.36), and low topic consistency (original: 0.5, normalized: -1.42), contributing to the model’s prediction of NCD with a high probability (0.99).
VI-D Text-Visual Alignment Differences between NCDs and HCs
Figure 7: Relative differences in text-visual correlation maps between NCD and HC groups: Cool colors represent areas of lower correlation in the NCD group; warm colors represent areas of higher correlation in the NCD group. The text timeline is resampled (via undersampling or interpolation) to match the number of images for visualization.
Fig. 7 presents the relative differences in cross-modal alignments between individuals of HCs and NCDs. To match the varying lengths of text segments with the number of images, the text timeline is resampled (through under-sampling or interpolation) to a fixed length of 15 for visualization. The colormap highlights regions of text-visual correlation divergence, with cool colors indicating diminished alignment in NCD narratives and warm colors representing higher correlation for the NCD group. Notably, the NCD narratives exhibit relatively lower text-visual correlation across the heatmap diagonal, signifying their disrupted cross-modal consistencies in the VSN task. These findings demonstrate TITAN’s ability to model cross-modal consistency for narrative analysis, providing a new dimension for NCD detection.

### v3

@@ IV-D TITAN: Text-Image Temporal Alignment Network
While DTM demonstrates potential in capturing topic evolution patterns, it faces challenges in incorporating cross-modal information in tasks such as picture description and VSNs. This limitation impairs topic representation modeling, preventing the effective utilization of image information and the handling of cross-modality correspondence. Previous studies have attempted to address these issues by inputting images into multimodal large language models to generate reference descriptions and keywords [41]. However, these approaches often overlook temporal dynamics and fail to capture topic evolution throughout the entire narrative. To overcome these challenges, and drawing inspiration from multimodal metrics like CLIPScore and GROOViST [83, 84] (which measure the text-image alignment through contrastive learning), we propose the Text-Image Temporal Alignment Network (TITAN). TITAN treats images as dynamic topics and constructs temporal relationships using the Attention mechanism, as illustrated in Fig. 2.
This network comprises several key components: (1) feature extraction, (2) embedding fusion, (3) projection layers, (4) Attention module, and (5) prediction head. We describe each component in detail below.
@@
The first step involves extracting features from the input text and images using the pretrained text-visual models, comprising textual-encoder and visual-encoder to process different modalities. In this work, we adopted the ViT visual-encoder and RoBERTa textual-encoder from the pretrained CLIP-like models, as mentioned in Section IV-B. The CLIP-like models can effectively align the embeddings of images and corresponding texts in the semantic space, thus providing insights for narrative coherence (structural dynamics) and thematic consistency (content alignment). Specifically, given an input image $I$ and text $T$, we obtain the embeddings $E_{I}$ (from the [CLS] token) and $E_{T}$ (via mean pooling) from the final projection layer of their respective encoders:
$E_{I}=\text{CLIP}_{\text{Visual}}(I)\in\mathbb{R}^{J\times H},E_{T}=\text{CLIP}_{\text{Textual}}(T)\in\mathbb{R}^{K\times H}$
where $H$ is the hidden dimension of the extracted embeddings, $J$ is the number of image frames (e.g., $J=1$ for the ADReSS and ADReSSo corpora and $J=15$ for the CU-MARVEL-RABBIT corpus), and $K$ is the temporal length of the investigated text sequence.
@@
where $[;]$ denotes the concatenation operation. This concatenation operation offers at least two advantages. First, it allows the subsequent layers to share the same projection layers for both image and textual embeddings, which is parameter-efficient and feasible since the embeddings are aligned to a common multimodal space during the pre-training stage. Second, inputting the concatenated embeddings into the Attention mechanism enables the model to capture combined information from image-image, image-text, and text-text pairs, thereby enhancing the richness of the contextual features.
@@
As shown in Fig. 2(b), the Attention module aims to capture the temporal dynamics and contextual relationships within and between modalities, by attending to specific positions of the projected embeddings.
To effectively process dynamic temporal information, we incorporate Rotary Position Embedding (RoPE) [85] into the Attention mechanism. RoPE encodes positional information into token embeddings without disrupting the invariance properties of Self-Attention, making it well-suited for capturing temporal dependencies. Mathematically, the positional encoding vector for the $i$-th position, $p_{i}\in\mathbb{R}^{2H^{\prime}}$, is defined using sinusoidal functions:
@@
Each element of $p_{i}$ is derived from sinusoidal functions to capture positional information: $\sin(i)$ and $\cos(i)$ are applied to alternate dimensions, with increasing frequencies across dimensions to capture information at various scales. Given the different temporal strides of images and texts, RoPE is applied separately to each modality with different frequencies, i.e., $p_{i}\in\mathbb{R}^{1,\ldots,H^{\prime}}$ for images and $p_{i}\in\mathbb{R}^{H^{\prime}+1,\ldots,2H^{\prime}}$ for texts.
@@
where $Q^{\prime}$ and $K^{\prime}$ are the RoPE-transformed queries and keys. The mask ensures that the padding positions of input get near-zero Attention weights after the softmax. The scaling factor $\sqrt{H^{\prime}}$ prevents the dot products from growing too large in magnitude. The values remain untransformed to preserve the original contents that are being passed through the network.
In VSN tasks such as the Rabbit Story, narrative transcripts must align with reference images. Participants with NCDs may misidentify animal names (anomia) or omit detailed scenarios (visual perception deficts) in these images. For example, many NCDs misidentify “goose” as “duck”, or skip the event of “goose in the brush” in the seventh image. The Attention module in the proposed text-image temporal alignment network is able to perceive potential misalignment and discrepancies cross modalities, and therefore recognize NCD participants with the consideration of their narrative macrostructure.
D5. Prediction Head
Finally, the up-projected embedding is temporally pooled with the Attention mask ($\in\mathbb{R}^{J+K}$) and passed through a fully connected layer to generate the prediction:
$\hat{y}=\sigma(E_{pool}W_{fc}^{T}+b_{fc})\in\mathbb{R}^{C}$
where $C$ is the prediction dimension (number of classes for classification tasks, or 1 for the regression task), $W_{fc}\in\mathbb{R}^{C\times H}$ and $b_{fc}\in\mathbb{R}^{C}$ are the weight and bias of the fully connected layer, and $\sigma$ is an activation function (softmax for classification tasks and identity for the regression task).
Throughout the network, residual connections are applied at each stage to facilitate gradient flow and mitigate the vanishing gradient problem.
The TITAN model leverages high-quality cross-modal representations from CLIP-like models as input features, which inherently capture strong image-text correlations. As shown in Fig. 2(c), the heatmap visualizes the pairwise similarity between text chunks (x-axis) and image sequences (y-axis) computed from the raw CLIP embeddings $E_{T}$ and $E_{I}$. The diagonal intensity pattern indicates that these input features inherently exhibit temporal and cross-modal consistency in narrative structures. Bright spots along the diagonal reflect strong alignment between images and corresponding text segments, suggesting that events or scenes unfold in a logically ordered sequence. These initial representations are further refined through the joint design of the projection, Attention, and prediction modules, enabling TITAN to capture cross-modal consistency and temporal coherence for NCD detection.
V Experiments & Results
TABLE III: Classification and Regression Results on the CU-MARVEL-RABBIT corpus. Feature dimensions are given in parentheses.
“*” denotes baseline systems; $\downarrow$” indicates that the lower values are better; “Emb.” refers to embedding(s).
The best results are highlighted in red, and the second-best are underlined.
@@ V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and models. Eight systems are designed, each utilizing a distinct feature set. System 1 uses 10 acoustic features, System 2 incorporates 13 linguistic features, and System 3 combines these microstructural features. System 4 utilizes 16 proposed reference-based metrics, and System 5 employs six proposed DTM-based features. System 6 combines these macrostructural features, while System 7 integrates all microstructural and macrostructural features. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with five components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and a soft margin $C=1$. The number of principal components in PCA and the SVM parameters (namely, the “kernel” and “soft margin”) are tuned via grid search using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 2, has a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as five.
During training, the models are optimized for 100 epochs with an initial learning rate of 1e-3 and a weight decay of 1e-2. The pretrained CLIP models are frozen, except the last projection layer of the textual encoder, which is fine-tuned with a learning rate of 1e-5. The batch size is set to 64 for the ADReSS and ADReSSo corpora and 256 for the CU-MARVEL-RABBIT corpus to accommodate differences in data availability and computational efficiency.
@@ V-B Evaluation Protocol
For evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, the primary metrics were the F1-score and AUC, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Except for AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we normalized the ground-truth labels to the range [0, 1] and used the coefficient of determination ($R^{2}$, ranging from -$\infty$ to 1) and Root Mean Square Error (RMSE, ranging from 0 to $\infty$) as metrics. To ensure robustness in evaluation, results were averaged over the five final epochs of each experiment. Notably, the testing sets of ADReSS, ADReSSo, and CU-MARVEL-RABBIT were held out from training and approximately balanced in terms of positive and negative samples, ensuring an unbiased evaluation.
@@ V-C Results
Table III presents the results of classification and regression tasks, comparing various feature sets and models using F1 as the primary evaluation metric. Among the systems, the proposed TITAN model (Sys. 8), which utilizes combined text and image embeddings, achieves the best overall performance in both tasks, with an F1 score of 0.7238 and AUC of 0.8120 in classification; and R2 of 0.3876 in regression. In contrast, the linguistic feature-based system (Sys. 2) shows an inferior performance among these systems, which is limited by the ASR, segmentation and tagging components that are not tailored to colloquial Cantonese narratives.
Another key finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6947 versus 0.5843 in classification; and R2 of 0.2233 versus 0.1604 in regression. This highlights the effectiveness of discourse-level analysis in capturing NCD-related narrative characteristics.
@@ VI Analysis & Discussion
In this section, we discuss the proposed DTM-based features and the detailed mechanisms of TITAN. We first present an instance-level case study to illustrate these DTM-based features. We then extend the analysis to all test examples to examine the correlation between the statistical features and NCDs, as well as their contributions to the classification model’s performance. Next, we investigate the interpretability of TITAN through the visualization of its Attention mechanism, and evaluate the impact of individual components via ablation studies. Finally, we contextualize our findings by comparing TITAN with prior research on the ADReSS and ADReSSo corpora.
Figure 3: Correlation coefficients between proposed features and NCD labels, with statistical significance shown on the x-axis. The heights of the green (macrostructural) and red (microstructural) bars represent correlation values, indicating that most top-ranked features are macrostructural.
VI-A Case Study
To facilitate an intuitive understanding of the DTM-based features introduced in Section IV-C, we present a comparative case study between a healthy control participant and a participant with NCDs from the CU-MARVEL-RABBIT corpus. The transcripts and normalized feature values for these two examples are shown in Figs. 6 and 7 in the Appendix.
As shown in Fig. 6, the HC participant produced a coherent and lexically diverse narrative. The transcript captures most salient visual elements, corresponding to a high “topic consistency” score (0.50). The mention of “back home” is reflected in a positive “topic cycle” value (normalized as 1.26). High “topic variability” (1.37) indicates rich lexical diversity within each topic distribution. The “topic ptp range” (1.12) reflects that specific topics are prominently expressed at each time step, i.e., topic distributions across time steps exhibit high peak-to-peak amplitudes. The negative “topic change rate” (-0.47) suggests smooth, gradual topic transitions. The strong “topic temporal correlation” (0.97) indicates that topic progression follows a coherent, time-ordered storyline.
In contrast, the NCD narrative (Fig. 7) omits or misidentifies key elements (e.g., incorrect animals such as cat, duck, chick, and sheep), resulting in low “topic consistency” (-0.78). The absence of narrative closure (“return home”) leads to a negative “topic cycle” value (-0.80). Low “topic variability” (-2.46) indicates limited vocabulary for each topic. A high “topic change rate” (0.56) and a negative “temporal correlation” (-1.61) indicate incoherent topic development.
@@ VI-B Statistical Features Correlation Ranking
Building upon the insights from the case study, we extend our analysis to assess how the proposed statistical features are associated with NCDs across all examples in the CU-MARVEL-RABBIT corpus. To this end, we conduct Spearman’s rank correlation test across features and ranked them by the absolute value of correlation coefficients. As illustrated in Fig. 3, the correlation distribution shows that the macrostructural features (green bars) exhibit greater importance compared to microstructural features (red bars). Notably, the DTM-based metric “topic consistency”, the acoustic feature “speaking rate”, and several reference-based coverage metrics ranked relatively high with strong statistical significance ($p<0.001$), suggesting their potential as indicators of NCDs. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features show positive correlations with NCDs. These findings highlight the utility of these features not only for effective classification, but also for gaining deeper insight into NCD-related narrative characteristics.
@@ VI-C Interpretable Feature Contribution Analysis
To elucidate the role of the proposed features in the model’s decision pathways, we quantify their contributions using SHapley Additive exPlanations (SHAP) values [86]. SHAP is a model-agnostic explanation framework rooted in cooperative game theory, which fairly and consistently attributes a model’s prediction to each input feature using Shapley values. Specifically, for each feature $x_{i}\in X$ (the complete feature set), the SHAP value $\phi_{i}$ represents its average marginal contribution to the model’s predictions when the target feature is added to all possible subsets $S$ of the other features, i.e.,
@@
where $w_{i}=|S|!(|X|-|S|-1)!/|X|!$ denotes the weight accounting for subset permutations, $|\cdot|$ denotes set cardinality, and $f(S)=ln{\frac{P(\hat{y}|S)}{1-P(\hat{y}|S)}}$ is the log-odds ratio of the model’s output probability $P$ based on given inputs $S$.
@@
This formulation enables global feature importance analysis by aggregating SHAP values over the held-out test set. Fig. 4 presents the global feature contributions for System 7, ranked by their absolute SHAP values. Notably, the proposed macrostructural features, including “topic variability”, “topic change rate”, “topic consistency” and “ROUGE-L”, emerged as the top-ranked global contributors. These findings highlight the critical role of macrostructural patterns in distinguishing between individuals with NCDs and HCs. The consistent directionality of their SHAP values (positive or negative associations with NCD classification) further validates their robust discriminative power across diverse instances.
Figure 4: Global SHAP value distribution (beeswarm plot) across all test samples in the CU-MARVEL-RABBIT corpus. Features are ranked vertically by mean absolute SHAP values (most impactful at the top). Each point represents an instance: the horizontal position indicates the SHAP value (negative or positive impact), the color reflects the standardized feature value (blue = low, red = high), and density of points along each feature indicates value distribution.
VI-D Attention Mechanism in Cross-modal Consistency
The Attention-based analysis illustrates how TITAN refines and leverages cross-modal consistency for NCD detection, as illustrated in Fig. 5. Fig. 5(a) shows the raw cross-modal alignment differences between NCD and HC groups using initial CLIP embeddings. The colormap highlights regions of text-visual correlation divergence, with warm colors indicating higher alignment and cool colors representing lower alignment for the HC group. Fig. 5(b) displays position-wise cosine similarities from RoPE embeddings, revealing an inherent structure that encodes temporal proximity.
It should be noted that although TITAN operates on raw textual and visual embeddings rather than their correlation maps, the learned Attention weights as shown in Fig. 5(c) and (d) can effectively focus on discriminative text-visual relationships. Fig. 5(c) shows that for the HC group, Attention highlights temporally coherent regions consistent with RoPE similarities. Fig. 5(d) shows that for the NCD group, Attention shifts towards misaligned or inconsistent cross-modal parts. This adaptive weighting enhances the model’s sensitivity to cognitive impairments and thereby improves the performance of NCD detection.
Figure 5: Attention mechanism in text-visual alignment. (a) Raw text-visual correlation differences between the NCD and HC groups. (b) Raw position-wise similarities of RoPE. Text-visual attention weights for (c) the HC group (c) and (d) the NCD group, with texts as the query and images as the key. For visualization purposes, the text timeline is resampled (via downsampling or interpolation) to match the number of images.